Suppose that 10 years ago you bought a home for $120,000, paying 10% as a down payment, and financing the rest at 9% interest for 30 years.

How much interest have you paid so far (over the last 10 years)?
1
Expert's answer
2018-09-12T09:35:08-0400
Suppose that 10 years ago you bought a home for $120,000, paying 10% ($12,000) as a down payment,
and financing the rest at 9% interest for 30 years.
In this case total sum to pay is:
FV = (120,000 - 12,000)*1.09^30 = 108,000*1.09^30= $1,432,909.27.
In 10 years (10 years is 1/3 of 30 years) you will pay 1,432,909.27/3 = $477,636.43,
where 108,000/3 = $36,000 is a part of the loan repaid and $477,636.43 - $36,000 = $441,636.43 is interest paid over 10 years.